# Exotic compact objects and light bosonic fields

### v1

@@ Abstract
In this note, we discuss the effect of light, non-gauge, bosonic degrees of freedom on the exterior spacetime of an exotic compact object. We show that such fields generically introduce large deviations from spacetimes of vacuum General Relativity near and outside the surfaces of ultra-compact exotic objects unless one assumes they totally decouple from the standard model or new heavy fields. Hence, using solutions of vacuum General Relativity to model ultra-compact exotic objects and their perturbations relies implicitly on this assumption or on the absence of such fields.
@@
where $M$ is the ADM mass of the ECO’s spacetime, $Q$ is the scalar charge, that is the $1/r$ coefficient that appears in the expansion of the scalar field at infinity as a power series in $1/r$ i.e., $\phi(r\rightarrow\infty)=\phi_{\infty}+\frac{Q}{r}+\mathcal{O}\left(1/r^{2}\right)$. For the perturbative equations and the relation between $M,Q,m$, and $c$ see appendix A.
@@
Our arguments do not rely on the nature or specific interactions of the light bosonic fields, nor do they assume that they are crucial for the prevention of collapse and the formation of the horizon - they merely have to be present. This is because if they exist, their configurations tend to diverge near the would-be horizon unless the ECO carries zero charges associated with these fields. The fact that we have not yet detected new light fields in non-gravitational experiments does suggest that, if they exist, they couple weakly to the standard model at low curvatures. However, to assume that the ECO carries zero charge translates to a much stronger assumption about the coupling of these light fields, the standard model fields, and new heavy fields in the interior of the ECO.
For simplicity, we have used spherical symmetry in our analysis. Relaxing this assumption does not remove the issue we point out regarding light fields but instead introduces an additional complication. For less symmetric spacetimes the properties of the interior are encoded in the multipolar structure of the exterior. Axisymmetric spacetime is a characteristic example, where one has a plethora of solutions with different properties already in general relativity [48, 49, 50, 51, 52]. Hence, even if Eq. (1) would act as a good approximation for the exterior, it would be very tenuous to suggest that any particular solution of these equations can generically act as an adequate approximation for the spacetime of an ECO near its surface.
@@ Acknowledgement
We would like to thank Andrea Maselli and Silke Weinfurtner for useful discussions. G.V. acknowledges support from the Czech Academy of Sciences under grant number LQ100102101. T.P.S. acknowledges partial support from the STFC Consolidated Grants no. ST/X000672/1 and no. ST/V005596/1.

### v3

@@ Abstract
In this note, we discuss the effect of light, non-gauge, bosonic degrees of freedom on the exterior spacetime of an exotic compact object. We show that such fields generically introduce large deviations from black hole spacetimes of General Relativity near and outside the surfaces of ultra-compact exotic objects unless one assumes they totally decouple from the standard model or new heavy fields. Hence, using black hole spacetimes of General Relativity to model ultra-compact exotic objects and their perturbations relies implicitly on this assumption or on the absence of such fields.
@@
where $M$ is the ADM mass of the ECO’s spacetime, $Q$ is the scalar charge, that is the $1/r$ coefficient that appears in the expansion of the scalar field at infinity as a power series in $1/r$ i.e., $\phi(r\rightarrow\infty)=\phi_{\infty}+\frac{Q}{r}+\mathcal{O}\left(1/r^{2}\right)$. For the perturbative equations and the relation between $M,Q,m$, and $c$ see appendix A. It is worth noting that eqs. (5) and (6) are known to admit the Janis-Newman-Winicour metric as an exact solution [45, 46]. Our perturbative solution agrees to second order in $Q$ with a small charge expansion of this exact solution. We restrict ourselves to perturbative treatment, as it is more transparent and easy to follow.
@@
Our arguments do not rely on the nature or specific interactions of the light bosonic fields, nor do they assume that they are crucial for the prevention of collapse and the formation of the horizon - they merely have to be present. This is because if they exist, their configurations tend to diverge near the would-be horizon unless the ECO carries zero charges associated with these fields. The fact that we have not yet detected new light fields in non-gravitational experiments does suggest that, if they exist, they couple weakly to the standard model at low curvatures. However, to assume that the ECO carries zero charge translates to a much stronger assumption about the coupling of these light fields, the standard model fields, and new heavy fields in the interior of the ECO. A possible workaround would be to devise a screening mechanism where the nonlinearities in the system conceal the scalar field all the way to the surface of the smallest ECO. Doing so would allow for a non-perturbative restoration of GR outside the ECO.
For simplicity, we have used spherical symmetry in our analysis. Relaxing this assumption and assuming that the exterior is described by the Kerr metric does not remove the issue we point out regarding light fields. For other vacuum spacetimes of general relativity the properties of the interior are encoded in the multipolar structure of the exterior. Axisymmetric spacetimes are a characteristic example, where one has a plethora of solutions with different properties already in general relativity [50, 51, 52, 53, 54]. Hence, even if Eq. (1) would act as a good approximation for the exterior, it would be very tenuous to suggest that any particular solution of these equations can generically act as an adequate approximation for the spacetime of an ECO near its surface.
@@ Acknowledgement
We would like to thank Pedro Fernandes, Andrea Maselli and Silke Weinfurtner for useful feedback on earlier versions of this manuscript. G.V. acknowledges support from the Czech Academy of Sciences under grant number LQ100102101. T.P.S. acknowledges partial support from the STFC Consolidated Grants no. ST/X000672/1 and no. ST/V005596/1.